OpenStudy (anonymous):

Solve the system of equations using the elimination method. 3x - 2y = -8 2x - 3y = -7

OpenStudy (anonymous):

you want to make x or y the same so lets make x the same multiply all of the first equation by 2 and second equation by 3

OpenStudy (anonymous):

6x-4y=-16 6x-9y=-21

OpenStudy (anonymous):

6x cancel out

OpenStudy (anonymous):

@bahrom7893 Give me help

OpenStudy (anonymous):

im lost

OpenStudy (bahrom7893):

eliminate a variable: 3x - 2y = -8

OpenStudy (anonymous):

So far so good

OpenStudy (bahrom7893):

3x=-8+2y x=(-8+2y)/3 plug that into the 2nd equation: 2(-8+2y)/3 - 3y = -7

OpenStudy (bahrom7893):

solve for y and then plug that back into the first equation

OpenStudy (anonymous):

is the answer no Solution?

OpenStudy (anonymous):

algebra look at bahrom's masterpeice there is a solution, we cant give you the answers but we have told you how to do it

OpenStudy (anonymous):

masterpiece*

OpenStudy (anonymous):

(-1,-2) so this is the answer i got is it correct?